OpenStudy (anonymous):

The function below shows the number of students of a school who enrolled for cooking classes. Let f(x) represent the total number of students who enrolled for the classes after x years. f(x) = 11(1.35)x The average rate of change in the number of students who enrolled for cooking classes from the first to the fifth year is ________students per year. Round your answer to the nearest whole number.

OpenStudy (anonymous):

Average rate of change over an interval \([a,b]\) is given by \[\frac{f(b)-f(a)}{b-a}\] Here, \([1,5]\) is your interval.

OpenStudy (anonymous):

i dont quite understand

OpenStudy (anonymous):

What is \(f(x)\)? What is \(f(b)=f(5)\)? And what is \(f(a)=f(1)\)?

OpenStudy (anonymous):

Plug in the end points of the interval, evaluate the function at those endpoints. The average rate of change is the ratio of (the difference of the function evaluated at the endpoints) to (the difference of the values of the endpoints).